Robin Design, Inc. is a Web site design and consulting firm. The firm uses a job order costing system in which each client isRobin Design, Inc. is a Web site design and consulting firm. The firm uses a job order costing system in which each client isRobin Design, Inc. is a Web site design and consulting firm. The firm uses a job order costing system in which each client isRequirement 2. Compute the total cost of each job. First, enter the direct costs for each job. Then enter in the indirect cosRequirement 4. Why does Robin Design assign costs to jobs? Robin Design assigns costs to jobs to help the company that cover
Answer #1

Robin Design Inc

  1. Computation of the direct labor rate and predetermined overhead allocation rate for 2018:

Direct labor rate –

Estimated Direct Labor costs

Estimated direct labor hours

=

Direct labor cost

24,00,000

8,000

=

300

Predetermined overhead allocation rate for 2018 –

Estimated overhead costs –

Support staff salaries        $354,000

Computer leases               $44,000

Office Supplies                 $22,000

Office Rent                       $60,000

Estimated OH Costs         $480,000

Allocation rate = (estimated overhead cost/direct labor cost) x 100

= 480,000/2,400,000 = 20%

  1. Computation of total cost of each job:

Robin Design Inc

Estimated Cost of Tasty Coop and Maynard Chocolates Jobs

Tasty Coop

Maynard Chocolates

Direct Costs:

Direct labor

800 hours x $300

$240,000

500 hours x $300

$150,000

Software licensing costs

$2,500

$500

Travel Costs

$7,000

0

Total Direct Costs

$249,500

$150,500

Allocated indirect costs:

20% x 240,000

$48,000

20% x 150,000

$30,000

Total Costs

$297,500

$180,500

  1. Desired profit = 20% of service revenue

Service revenue = ?

Service revenue = cost + profit

Cost = service revenue – profit

If service revenue = 100%

Profit = 20%

Then cost = 80% of service revenue

Applying these percent to each of the jobs –

Tasty Coop –

Cost = 297,500

Service revenue = cost/80%

Service revenue = 297,500/80% = $371,875

Maynard Chocolates –

Cost = $180,500

Service revenue = cost/80%

Service revenue = 180,500/80% = $225,625

Service Revenue

Tasty Coop = $371,875

Maynard Chocolates = $225,625

  1. Why assign costs to jobs:

Robin Design assigns costs to jobs to help the company estimate service revenue that cover all costs and contribute to profit. Assigning costs to Jobs also can help Robin Design control costs.